OpenStudy (anonymous):

Every graduating senior from Dr. John Hugh Gillis High enters St. Francis Xavier University, as do one half of those from Guysborough High. At St. F.X., one half of the freshman class are from Dr. John Hugh Gillis High. If 400 graduates from Guysborough High and 300 from Dr. John Hugh Gillis, how many of St.F.X. freshmen are neither from Dr. J. H. Gillis nor Guysborough High?

OpenStudy (dumbcow):

300 represents half the freshman class because all 300 graduates from Dr.Johns go to university. Half of Guysborough graduates go so thats 400/2 = 200 Total amount of freshman is twice 300 -> 300*2 = 600 600 - 300-200 = 100 100 Freshman are neither from DrJohns or Guysborough